I have tried this but I am having trouble, can you please help. Use substitution to solve each system. If it does not have one solution then put no solution or indinitely many solutions. 1/3X-Y=3 2X+Y=25
Answer by checkley77(12569)

1/3X-Y=3
-y=-x/3+3
y=x/3-3 now replace y with (x/3-3) in the other equation
2X+Y=25 SORY I MISSED THE 2 IN FRONT O THE X.
2x+(x/3-3)=25
2x+x/3=25+3
2x+x/3=28
(2x*3+x)/3=28
(6x+x)/3=28
7x=3*28
7x=84
x=84/7
x=12 ans.
y=12/3-3
y=4-3
y=1 ans.
Proof:
2*12+1=25
24+1=25
25=25